# The Internal Model Principle of Time-Varying Optimization

### v1

@@ 4 $F_{c}(z,y)\leftarrow s(z)+L(y-\nabla_{x}f(H_{c}(z),z))$;
Remark 5.
Instead of a Luenberger observer, alternative dynamic observers could be considered in Line 4 of the algorithm to achieve different asymptotic or transient properties of the resulting gradient-feedback optimization algorithm; we leave the investigation of alternative state observer algorithms as the scope of future works. $\square$
We illustrate the applicability of (13) on a quadratic problem in the following example.
@@
Consider the quadratic problem studied in Example 3, and assume that the exosystem follows the linear model $\dot{\theta}=S\theta$ for some matrix $S\in{}^{p\times p}.$ According to Theorem 5, an optimization algorithm given by
@@
00 0 $=(RG_{c}\Sigma+Q),$ (18b)
for all $\theta$ in a neighborhood of the origin. When this condition holds, the application of Algorithm 1 gives
$A_{c}=S,$ $B_{c}=L,$ $G_{c}=-R^{\dagger}Q,$
where $L$ is any matrix such that $S-LQ$ is Hurwitz; notice that this choice satisfies (18) with $\Sigma=I.$square
IV-B The optimization algorithm as a copy of the exosystem
As discussed after (16), the state of any optimization algorithm that achieves asymptotic tracking and that of the exosystem must be related by a change of coordinates. In the following, we show that $\sigma(\theta)$ is indeed an injective map, and it can thus be interpreted as a change of coordinates. We begin with the following result.
Proposition 6.
@@ IV-B The optimization algorithm as a copy of the exosystem
The claim follows by noting that $F_{c}(z,y)$ and $G_{c}(z)$ obtained by Algorithm 1 satisfy (13) with $\sigma$ the identity operator on $\Theta$. ∎
The interpretation of Proposition 6 is that, for any solution returned by Algorithm 1, $F_{c}(z,y)$ incorporates a copy of $s(\theta);$ precisely,
$F_{c}(\sigma(\theta),0)=s(\theta).$ (19)
The above feature is not a result of using Algorithm 1. Indeed, every gradient feedback algorithm has this feature, as shown next.
Proposition 7.
@@
By contradiction, assume that $F_{c}(z,y)$ and $G_{c}(z)$ satisfy (13), but $\sigma(\theta)$ is not injective at the origin; namely, there exists nonzero $\theta^{\prime}\in\Theta_{\circ}$ such that $\sigma(\theta^{\prime})=z_{0}^{\star}.$ From (13b), we have
$0=\nabla_{x}f(G_{c}(\sigma(\theta^{\prime})),\theta^{\prime}).$
Moreover, from Theorem 4,
@@
For $C^{\infty}$ vector fields $h_{1}(x)$ and $h_{2}(x)$, we let
$L_{h_{1}}(h_{2})(x)=\frac{\partial h_{2}(x)}{\partial x}h_{1}(x).$
By application of (20), we have
@@
Hence, the matrix
@@
is not invertible at $\theta=\theta^{\prime}.$ By [35, Thm. 3.13], the system (2) is not weakly observable, thus violating Assumption 4. ∎
V Fidelity of the internal model and tracking accuracy
In this section, we illustrate through examples how an imprecise knowledge of the internal model impacts the tracking accuracy. For tractability, we will focus on the quadratic problem (12) (see Examples 3 and 4). Suppose that an imprecise knowledge of the internal model is available; namely, there exists $\Delta\in{}^{n_{c}\times p}$ such that (18) is modified to:
$\Sigma S+\Delta$ $=A_{c}\Sigma,$ (21a)
00 0 $=(RG_{c}\Sigma+Q).$ (21b)
@@ V Fidelity of the internal model and tracking accuracy
$\dot{\tilde{z}}=(A_{c}+B_{c}RG_{c})\tilde{z}+\Delta\theta.$ (22a)
The corresponding gradient signal in terms of $\tilde{z}$ is:
$y=RG_{c}\tilde{z}.$ (22b)
Assuming that $(A_{c}+B_{c}RG_{c})$ is Hurwitz stable (see Theorem 4), the Final Value Theorem gives:
@@
As a first illustrative scenario, suppose $S=0;$ namely, that the exosystem states are constants at all times. Notice that, since $S=0,$ from (18a), $A_{c}$ incorporates an internal model of $S$ if and only if $A_{c}\Sigma=0;$ precisely, if and only if
$\text{dim}(\ker(A_{c}))\geq p.$
In this case, assuming that the perturbed $A_{c}$ is such that $(A_{c}+B_{c}RG_{c})$ remains Hurwitz stable, from (23), we have:
@@
As expected, when $\Delta=0,$ $A_{c}$ incorporates an exact internal model, and $y_{\infty}=0$ for any $\theta(0).$ On the other hand, when $\Delta\neq 0,$
@@
Interestingly, analogous continuity properties may not hold when the Poisson stability assumption is dropped. Consider an instance of (12) with $n=p=2,$ $R=Q=I,$ and
$S=\begin{bmatrix}0&1\\ 0&0\end{bmatrix}.$
Suppose the algorithm (17) is used, with $G_{c}=I,B_{c}=I,$
$A_{c}=\begin{bmatrix}-\epsilon_{1}&1\\ 0&-\epsilon_{2}\end{bmatrix},$
where $0\leq\epsilon_{1}<1,$ $10\leq\epsilon_{2}<$ 1 . To avoid trivial cases, we will assume $\theta_{1}(0)$ and $\theta_{2}(0)$ are nonzero. When $\epsilon_{1}=\epsilon_{2}=0,$ (21) holds with $\Sigma=-I$ and $\Delta=0$, in which case $y_{\infty}=0.$ On the other hand, when $\epsilon_{1},\epsilon_{2}\neq 0,$ from (23):
$y_{\infty}=\lim_{s\rightarrow 0}\begin{bmatrix}\frac{\epsilon_{1}}{s-\alpha_{1}}&\frac{\epsilon_{1}}{s(s-\alpha_{1})}+\frac{\epsilon_{2}}{(s-\alpha_{1})(s-\alpha_{2})}\\ 0&\frac{\epsilon_{2}}{s-\alpha_{2}},\end{bmatrix}\begin{bmatrix}\theta_{1}(0)\\ \theta_{2}(0)\end{bmatrix}$
where $\alpha_{i}=1-\epsilon_{i}$ for $i\in\{1,2\}.$ Since $\epsilon_{1}\neq 0$ and $\theta_{2}(0)\neq 0$, we conclude that $|y(t)|\rightarrow\infty$ as $t\rightarrow\infty$, i.e., asymptotic tracking is not attained.
In conclusion, the tracking accuracy depends on the fidelity of the internal model as well as the asymptotic behavior of the exosystem.
VI Extensions to constrained problems
We can extend our results to study constrained time-varying optimization problems by searching for a stationary point of the Lagrangian.
Consider the equality-constrained problem
@@ VI Extensions to constrained problems
subject to $h_{i}(x,\theta(t))=0,\quad i=1,\ldots,m$
@@
$L(x,\lambda,\theta(t))=f(x,\theta(t))+\sum_{i=1}^{m}\lambda_{i}h_{i}(x,\theta(t))$
@@
$L(x,\bar{\lambda},\theta(t))\leq L(x,\lambda,\theta(t))\leq L(\bar{x},\lambda,\theta(t))$
for all pairs $(\bar{x},\bar{\lambda})\in{}^{n}\times{}^{m}$. For any such saddle-point, $x$ is primal optimal, $\lambda$ is dual optimal, and the optimal duality gap is zero. Moreover, the gradient of the Lagrangian (assuming it exists) is zero at any saddle-point. It follows from the derivations in the previous sections that the gradient-feedback and parameter-feedback algorithms can be directly applied to seek a stationary point of the Lagrangian function by replacing (in (1)) the variable $x$ with the extended decision variable $\tilde{x}=(x,\lambda)$ and by letting $f(\tilde{x},\theta)=L(x,\lambda,\theta).$ Notice that, if the critical point computed by (4a) is also a saddle-point, then it is also a solution to the equality-constrained problem; see [36, Ch. 5].
VII Simulation results
In this section, we illustrate our results and optimization design method through a set of numerical simulations.
Figure 3: Simulation results illustrating the performance of an optimization algorithm synthesized using Algorithm 1 for the quadratic instance (12) of (1). See Example 3 and Section VII-A for a discussion. (Top) Illustration of the temporal variability of the parameter $\theta(t)$ and of $z(t).$ (Second from top) $z(t)$ is an estimator for $\theta(t),$ and thus $z(t)\rightarrow\theta(t)$ as $t\rightarrow\infty.$ (Third from top) The proposed control algorithm is successful in regulating the gradient feedback signal $y(t)$ to zero asymptotically. (Bottom) Illustration that $x(t)\rightarrow x^{\star}(t)$ as $t\rightarrow\infty.$
VII-A Optimization design for quadratic costs
@@
$F_{c}(z,y)$ $=s(z)+L(y-\nabla_{x}f(H_{c}(z),z)),$
$=Sz+L(y-(RG_{c}z+Qz)),$
$=Sz+Ly.$
The simulation results are presented in Figure 3. It follows from our choice of $F_{c}(z,y),$ that the optimization state $z(t)$ converges to the parameter vector $\theta(t)$ (see the top two plots in Fig. 3). Moreover, it follows from our choice of $G_{c}(z)$ that $y(t)\rightarrow 0$ and thus $x(t)\rightarrow x^{\star}(t)$ as $t\rightarrow\infty$ (see the bottom two plots in Fig. 3).
@@
VII-B Application to solve the dynamic traffic assignment problem in transportation
We next illustrate the applicability of the framework in solving the dynamic traffic assignment problem in roadway transportation [37]; intuitively, the objective is to decide how traffic flows split among the available paths of a network to minimize the drivers’ travel time to destination. We model a roadway transportation network using a static flow model [37], described by a directed graph $\mathcal{G}=(\mathcal{V},\mathcal{E}),$ with edges $i\in\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$ (modeling traffic roads) and nodes $\mathcal{V}$ (modeling traffic junctions). For $i\in\mathcal{E},$ we denote by $i^{+}\in\mathcal{V}$ and $i^{-}\in\mathcal{V}$ its origin and destination nodes, respectively. We assume that an exogenous, time-varying, inflow of traffic $\theta(t)$ enters the network at a certain origin node, denoted by $o\in\mathcal{V},$ and exits at a certain destination node, denoted by $d\in\mathcal{V};$ for simplicity, we assume that there is only one origin-destination pair, but this is without loss of generality [37]. We describe the network state using a vector $x\in\mathbb{R}_{\geq 0}^{|\mathcal{E}|}$ (where $|\mathcal{E}|$ denotes the number of edges) whose entries $x_{i}$ describe the amount of inflow $\theta$ routed through road $i.$ To each link $i,$ we associate a function $\ell_{i}(x_{i})$ describing the latency (or travel time) of road $i$. For our simulations, we consider the network topology in Fig. 4, and choose the latency functions as follows:
@@ VII-B Application to solve the dynamic traffic assignment problem in transportation
According to Wardrop’s first principle [37, pp. 31], transportation networks operate at a condition where travelers select their path to minimize their travel time to destination. Mathematically, a Wardrop’s equilibrium is the optimizer of the following optimization problem:
@@
where $\delta_{v}(\theta),$ $v\in\mathcal{V},$ is defined as:
$\delta_{v}(\theta)=\begin{cases}\theta,&\text{if }v=o,\\ -\theta,&\text{if }v=d,\\ 0,&\text{otherwise}.\end{cases}$
The loss function in (VII-B) is used to model travelers who will switch to a different path if it has shorter travel time to destination, while the first constraint in (VII-B) describes the network topology, namely, that traffic flows are conserved at each node. Notice that (VII-B) is a time-varying optimization problem, where the temporal variability originates from the dependence of the constraint on $\theta(t)$, which describes the inflow of vehicles at the origin and outflow at the destination, measured in vehicles per hour. For our simulations, we assume that the network inflow is sinusoidal:
@@
where $\theta_{0},\theta_{1},\theta_{2},\omega_{1},\omega_{2},\phi_{1},\phi_{2}\in\mathbb{R}_{>0},$ satisfy $\theta_{0}>\theta_{1},$ $\theta_{0}>\theta_{2},$ and $\omega_{2}>\omega_{1}.$ The model (25) states that the network inflow is the sum of a constant term, $\theta_{0},$ a slowly-varying sinusoid with angular frequency $\omega_{1}$ and a quickly-varying sinusoid with angular frequency $\omega_{2}.$ The low-frequency sinusoid is used here to describe slowly-varying (e.g., hourly) traffic demands, while the high-frequency sinusoid is used to model sudden (e.g., at the minute-level) variations in traffic demand. For our simulations, we let $\theta_{0}=3$ veh/h, $\theta_{1}=1$ veh/h, $\theta_{2}=0.1$ veh/h, $\omega_{1}=0.1$ rad/hour, $\omega_{2}=\sqrt{50}$ rad/hour, and $\phi_{1}=\phi_{2}=0,$ see Fig. 5 (top).
We applied Algorithm 1 to derive an optimization algorithm to solve the traffic assignment problem (VII-B). For the synthesis, we utilized the internal model $s(z)=Sz,$ where
$S$ $=\begin{bmatrix}0&1&0&0&0\\ -\omega_{1}^{2}&0&0&0&0\\ 0&0&0&1&0\\ 0&0&-\omega_{2}^{2}&0&0\\ 0&0&0&0&0\end{bmatrix}.$
Notice that knowledge of $\theta_{0},\theta_{1},\theta_{2},\phi_{1},\phi_{2}$ is not required to synthesize the optimization algorithm – only the frequencies $\omega_{1},\omega_{2}$ are required to be known. To seek a solution to the constrained problem (VII-B), consider the Lagrangian function:

### v2

@@ 4 $F_{c}(z,y)\leftarrow s(z)+L(y-\nabla_{x}f(H_{c}(z),z))$;
Theorem 6 (Correctness of Algorithm 1).
Suppose Assumptions 13 hold. Then, the control algorithm (10) designed using Algorithm 1 asymptotically tracks a critical trajectory of (1).
Remark 13 (Alternative observer choices).
Instead of a Luenberger observer, alternative dynamic observers could be considered in Line 4 of the algorithm to achieve different asymptotic or transient properties of the resulting gradient-feedback optimization algorithm; we leave the investigation of alternative state observer algorithms to future works. $\square$
We illustrate the algorithm design procedure on a quadratic problem in the following example.
@@
Consider the quadratic problem studied in Example 2, and assume that the exosystem follows the linear model $\dot{\theta}=S\theta$ for some matrix $S\in{}^{p\times p}$ as in Assumption 2. Further, suppose that $Q$ is such that the unobservable subspace $\mathcal{N}=\cap_{i\geq 0}\ker(QS^{i})$ is the empty set $\mathcal{N}=\emptyset,$ so that Assumption 3 is satisfied; notice that this assumption holds, for example, when $n=p$ and $Q$ is invertible. According to Theorem 5, an optimization algorithm given by
@@
00 0 $=RG_{c}\Sigma+Q.$ (30b)
Next, we apply Algorithm 1 to design an algorithm that solves this problem. The mapping $H_{c}(\theta)=-R^{\dagger}Q\theta$ zeros the gradient (see Example 2). Following Algorithm 1, we set $n_{c}=p$, choose matrix $L$ such that $S-LQ$ is Hurwitz, define $G_{c}(z)=H_{c}(z)=-R^{\dagger}Qz$, and then select $A_{c}$ and $B_{c}$ such that $F_{c}(z,y)=A_{c}z+B_{c}y$ satisfies:
$F_{c}(z,y)$ $=s(z)+L(y-\nabla_{x}f(H_{c}(z),z)),$
$=Sz+L(y-(RH_{c}(z)+Qz)),$
$=Sz+Ly,$
which gives $A_{c}=S$ and $B_{c}=L.$ In summary, a dynamic gradient-feedback algorithm is (29) with $A_{c}=S$, $B_{c}=L$, and $G_{c}=-R^{\dagger}Q$, where $L$ is any matrix such that $S-LQ$ is Hurwitz. With this choice, conditions (30) are satisfied with $\Sigma=I$ (see Remark 12). Finally, the autonomous model (22) becomes $\dot{z}(t)=(S-LQ)z(t)$, which is asymptotically stable, as required by Theorem 4. $\square$
4.3 The optimization algorithm as a copy of the exosystem
As shown by (27), the state of optimization algorithms that achieve asymptotic tracking and that of the exosystem are related by a change of coordinates. In this section, we show that $\sigma(\theta)$ is an injective map and thus is a well-posed change of coordinates. For Algorithm 1, we have the following.
Proposition 7 (Injectivity of $\sigma(\theta)$ from Algorithm 1).
@@ 4.3 The optimization algorithm as a copy of the exosystem
The claim follows by noting that $F_{c}(z,y)$ and $G_{c}(z)$ obtained by Algorithm 1 satisfy (23) with $\sigma$ the identity. ∎
It follows from Proposition 7 that, for any optimization method returned by Algorithm 1, there exists an injective mapping that, locally, maps the state of the exosystem into the state of the optimization algorithm.
This property is not exclusive to Algorithm 1; in fact, any algorithm that achieves exact asymptotic tracking possesses the same characteristic, as demonstrated next.
Proposition 8 (Injectivity of any $\sigma(\theta)$).
@@
By contradiction, assume that $F_{c}(z,y)$ and $G_{c}(z)$ satisfy (23), but $\sigma(\theta)$ is not injective at the origin; namely, there exists nonzero $\theta^{\prime}\in\Theta_{\circ}$ such that $\sigma(\theta^{\prime})=z_{0}^{\star}.$ Then from (23b), $0=\nabla_{x}f(G_{c}(\sigma(\theta^{\prime})),\theta^{\prime})$. Moreover, from Theorem 4,
@@
For $C^{\infty}$ vector fields $h_{1}(x)$ and $h_{2}(x)$, we let $L_{h_{1}}(h_{2})(x)=\frac{\partial h_{2}(x)}{\partial x}h_{1}(x).$ By application of (31), we have
@@
Hence, the matrix of Lie derivatives of the measurable signal, which, intuitively, characterizes the information content of the gradient signal:
@@
is not invertible at $\theta=\theta^{\prime}.$ By [53, Thm. 3.13], the system (2) is not weakly observable, thus violating Assumption 3. ∎
5 Fidelity of the internal model vs accuracy
In Section 4, we have shown that a necessary condition for exact asymptotic tracking is that the algorithm incorporates an internal model of the temporal variability of the problem. Departing from this finding, in this section we tackle the following question: how will the algorithm perform when the temporal variability of the problem is known only approximately? For illustration purposes, we focus on the quadratic problem (19) (see Examples 2 and 3). Suppose that an imprecise internal model is available; that is, there exists $\Delta\in{}^{n_{c}\times p}$ such that (30) is modified to777 To explicitly model uncertainty in the matrix $S,$ equation (32) could be rewritten as $\Sigma(S+\Delta_{s})=A_{c}\Sigma.$ By comparison, the two approaches are equivalent with the substitution $\Delta=\Sigma\Delta_{s}.$ :
$\Sigma S+\Delta=A_{c}\Sigma\qquad\text{and}\qquad 0=RG_{c}\Sigma+Q.$ (32)
@@ 5 Fidelity of the internal model vs accuracy
$\dot{\tilde{z}}=(A_{c}+B_{c}RG_{c})\tilde{z}+\Delta\theta.$ (33)
The corresponding gradient signal in terms of $\tilde{z}$ is $y=RG_{c}\tilde{z}$. Assuming that $(A_{c}+B_{c}RG_{c})$ is Hurwitz stable (see Theorem 4), the Final Value Theorem gives:
@@
Interestingly, analogous continuity properties may not hold when the trajectories of (2) are not bounded (thus violating Assumption 2). Consider an instance of (19) with $n=p=2,$ $R=Q=I,$ and $S=\left[\begin{smallmatrix}0&1\\ 0&0\end{smallmatrix}\right]$. The solutions of (2) with this exosystem are a linear combination of modes $e^{0t}$ and $te^{0t},$ and thus grow unbounded at a linear rate. Suppose algorithm (29) is used with $G_{c}=I$, $B_{c}=I,$ and $A_{c}=\left[\begin{smallmatrix}-\epsilon_{1}&1\\ 0&-\epsilon_{2}\end{smallmatrix}\right]$, where $\epsilon_{1},\epsilon_{2}\in[0,1)$. When $\epsilon_{1}=\epsilon_{2}=0,$ (32) holds with $\Sigma=-I$ and $\Delta=0$, in which case $y_{\infty}=0$ and exact asymptotic tracking is achieved. On the other hand, when $\epsilon_{1},\epsilon_{2}\neq 0,$ (32) holds with $\Delta=\left[\begin{smallmatrix}\epsilon_{1}&0\\ 0&\epsilon_{2}\end{smallmatrix}\right]$ and, from (34), we obtain:
$y_{\infty}=\lim_{s\rightarrow 0}\begin{bmatrix}\frac{\epsilon_{1}}{s-\alpha_{1}}&\frac{\epsilon_{1}}{s(s-\alpha_{1})}+\frac{\epsilon_{2}}{(s-\alpha_{1})(s-\alpha_{2})}\\ 0&\frac{\epsilon_{2}}{s-\alpha_{2}}\end{bmatrix}\begin{bmatrix}\theta_{1}(0)\\ \theta_{2}(0)\end{bmatrix},$
where $\alpha_{i}=1-\epsilon_{i}$, $i\in\{1,2\}.$ Since $\epsilon_{1}\neq 0$, and provided that $\theta_{2}(0)\neq 0$, the above identity implies that $\|y(t)\|\rightarrow\infty$ as $t\rightarrow\infty$, i.e., the asymptotic tracking error grows unbounded.
This example illustrates that the tracking accuracy depends on the fidelity of the internal model but also the asymptotic behavior of the exosystem.
6 Extensions
In this section, we discuss possible extensions to constrained optimization problems and to discrete-time algorithms.
6.1 Constrained problems
Our results can be extended to constrained time-varying optimization. Consider the equality-constrained problem
@@ 6.1 Constrained problems
subject to $h_{i}(x,\theta(t))=0,\quad i=1,\ldots,m,$
@@
$L(x,\lambda,\theta(t))=f(x,\theta(t))+\sum_{i=1}^{m}\lambda_{i}h_{i}(x,\theta(t)),$
@@
$L(x,\bar{\lambda},\theta(t))\leq L(x,\lambda,\theta(t))\leq L(\bar{x},\lambda,\theta(t)).$
For any such saddle-point, if strong duality holds, $x$ is primal optimal, $\lambda$ is dual optimal, and the optimal duality gap is zero. Moreover, the gradient of the Lagrangian (assuming it exists) is zero at any saddle-point. It follows from the derivations in the previous sections that the gradient-feedback and parameter-feedback algorithms can be directly applied to seek a stationary point of the Lagrangian function by replacing (in (1)) the variable $x$ with the extended decision variable $\tilde{x}=(x,\lambda)$ and by letting $f(\tilde{x},\theta)=L(x,\lambda,\theta).$ Notice that, if the critical point computed by (10b) is also a saddle-point and strong duality holds, then it is also a solution to the equality-constrained problem. When strong duality does not hold, however, such a saddle-point may not correspond to an optimizer [54, Ch. 5].
6.2 Discrete-time algorithms
In practice, optimization algorithms are often implemented on digital machines and thus, in these cases, implementing (10) requires discretizations. In these cases, a possible solution is to pursue a design by emulation [55], which consists in converting the continuous-time algorithm (10) into a discrete-time one by approximation (that is, by approximating the time-derivative by a discrete difference using, e.g., Backward/Forward Difference, Tustin, ZOH methods, etc.). Then, the continuous-time guarantees given here can be translated into discrete-time ones (using standard tools from emulation [55] or sampled-data systems [56]), provided that the sampling time is sufficiently small. In these cases, exact tracking established here will continue to hold at the discrete time instants, while tracking will be inexact in-between sampling times and the error can be bounded in terms of the sampling frequency using the explicit bounds in [55].
An alternative method that could be pursued is that of discretizing the temporal variability $\theta(t)$, thus leading to a discrete-time optimization problem as in (4). For this alternative formulation, the critical trajectories are discrete signals and, in this case, it would be natural to seek a fully-discrete alternative to (10) and pursue an analysis in discrete time. Such a design is beyond the scope of this work and will be left as the focus of future works.
7 Simulation results
In this section, we illustrate our results and optimization design method through a set of numerical simulations. For implementation on digital machines, all algorithms have been discretized by using the explicit Runge-Kutta-Fehlberg method of order $4(5)$ with adaptive stepsize [57].
Figure 3: Simulation results illustrating the performance of an optimization algorithm synthesized using Algorithm 1 for the quadratic instance (19) of (1). See Example 2 and Section 7.1 for a discussion. (Top) Illustration of the temporal variability of the parameter $\theta(t)$ and of $z(t).$ (Second from top) $z(t)$ is an estimator for $\theta(t),$ and thus $z(t)\rightarrow\theta(t)$ as $t\rightarrow\infty.$ (Third from top) The proposed control algorithm is successful in regulating the gradient feedback signal $y(t)$ to zero asymptotically. (Bottom) Illustration that $x(t)\rightarrow x^{\star}(t)$ as $t\rightarrow\infty.$
7.1 Optimization design for quadratic costs
@@
7.2 Nonlinear example
We now illustrate our algorithm design on a nonlinear time-varying optimization problem from [30, 34], and we compare the results with several prediction-correction algorithm variants. Consider minimizing the function:
$f(x,\theta)=\tfrac{1}{2}(x-\theta)^{2}+\kappa\log(1+\exp(\mu x)),$
with parameters $\kappa=7.5$ and $\mu=1.5$, where the time-varying parameter $\theta(t)$ satisfies the exosystem dynamics $\dot{\theta}=S\theta$ where $S=\left[\begin{smallmatrix}0&1\\ 1&0\end{smallmatrix}\right]$. We compare the error $\|x(t)-x^{\star}(t)\|$ over time for several algorithms in Fig. 4. The objective function satisfies the conditions in Theorem 3; for our simulations, we computed numerically a mapping that zeros the gradient locally. In Fig. 4(a), we show the error for our algorithm with observer pole locations $(-\eta,-2\eta)$ for $\eta\in\{1,5,20\}$ (shown in dark to light blue) along with the corresponding rate $e^{-\eta t}$ (dashed black). For comparison, we also show the error of the prediction-correction (PC) algorithm (8) along with two of its finite-time (FT) variants [34, Equations (9) and (20)] using the parameter values in [34, Section VI.A].
Instead of computing $H_{c}(\theta)$ numerically, we can use its linear approximation in (20). Doing so yields the results in Fig. 4(b) for various size exosystem initial conditions $\theta(0)=\zeta\theta_{0}$ for $\zeta\in\{1,0.1,0.01\}$. The linear approximation does not zero the gradient exactly, which results in a finite steady-state error whose size depends on the size of the limit set of the exosystem (which corresponds to the size of the initial conditions since the exosystem dynamics are sinusoidal).
(a) Error over time for the proposed dynamic gradient-feedback (DGF) algorithm with observer pole locations $(-\eta,-2\eta)$ for $\eta\in\{1,5,20\}$ (dark to light blue) along with the prediction-correction (PC) algorithm and two finite-time (FT) variants.
(b) Error over time for the proposed dynamic gradient-feedback algorithm using the linear approximation to $H_{c}(\theta)$ in (20) with different size initial conditions $\theta(0)=\zeta\theta_{0}$ for $\zeta\in\{1,0.1,0.01\}$. When the initial conditions are small, the exosystem state stays near the equilibrium, in which case $H_{c}$ approximately zeros the gradient on the limit set of the exosystem.
Figure 4: Dynamic gradient-feedback algorithm applied to the nonlinear time-varying optimization problem from Section 7.2.
7.3 Application to solve the dynamic traffic assignment problem in transportation
@@
Figure 6: Illustration of the performance of the optimization method synthesized using Algorithm 1 to solve the dynamic traffic assignment problem (7.3). See Section 7.3 for a discussion. The bottom figure shows that, from a zero initial condition, the algorithm is capable of computing a Wardrop equilibrium in about $1.5$ hours, and then of tracking this equilibrium. Notice that, a faster rate of convergence could be obtained by shifting the eigenvalues of $S-LQ.$
We next illustrate the applicability of the framework in solving the dynamic traffic assignment problem in roadway transportation [58]; intuitively, the objective is to decide how traffic flows are split among the available paths of a network to minimize the drivers’ travel time to destination. We model a roadway transportation network using a static flow model [58], described by a directed graph $\mathcal{G}=(\mathcal{V},\mathcal{E}),$ with edges $i\in\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$ (modeling traffic roads) and nodes $\mathcal{V}$ (modeling traffic junctions). For $i\in\mathcal{E},$ we denote by $i^{+}\in\mathcal{V}$ and $i^{-}\in\mathcal{V}$ its origin and destination nodes, respectively. We assume that an exogenous, time-varying, inflow of traffic $\theta(t)$ enters the network at a certain origin node, denoted by $o\in\mathcal{V},$ and exits at a certain destination node, denoted by $d\in\mathcal{V};$ for simplicity, we assume that there is only one origin-destination pair, but this is without loss of generality [^8] When the network has multiple origin-destination pairs, an optimization problem of the form (7.3) will be associated with each origin-destination pair. Because the optimization problems are independent from one another [58], each of them can be solved independently using the methods illustrated here.. We describe the network state using a vector $x\in\mathbb{R}_{\geq 0}^{|\mathcal{E}|}$ (where $|\mathcal{E}|$ denotes the number of edges) whose entries $x_{i}$ describe the amount of inflow $\theta$ routed through road $i.$ To each link $i,$ we associate a function $\ell_{i}(x_{i})$ describing the latency (or travel time) of the road $i$. For our simulations, we consider the network topology in Fig. 5, and choose the latency functions as follows:
@@
According to Wardrop’s first principle [58, pp. 31], transportation networks operate at a condition where travelers select their path to minimize their travel time to their destination. Mathematically, Wardrop’s equilibrium is the optimizer of the following optimization problem:
@@
where $\delta_{v}(\theta)$ for $v\in\mathcal{V}$ is defined as $\delta_{o}(\theta)=\theta$ for the origin node, $\delta_{d}(\theta)=-\theta$ for the destination node, and $\delta_{v}(\theta)=0$ for all other nodes. The loss function in (7.3) is used to model travelers who will switch to a different path if it has a shorter travel time to destination, while the first constraint in (7.3) describes the network topology, namely, that traffic flows are conserved at each node. Notice that (7.3) is a time-varying optimization problem, where the temporal variability originates from the dependence of the constraint on $\theta(t)$, which describes the inflow of vehicles at the origin and outflow at the destination, measured in vehicles per hour. For our simulations, we assume that the network inflow is sinusoidal:
@@
where $\theta_{0},\theta_{1},\theta_{2},\omega_{1},\omega_{2},\phi_{1},\phi_{2}\in\mathbb{R}_{>0},$ satisfy $\theta_{0}>\theta_{1},$ $\theta_{0}>\theta_{2},$ and $\omega_{2}>\omega_{1}.$ The model (36) states that the network inflow is the sum of a constant term, $\theta_{0},$ a slowly-varying sinusoid with angular frequency $\omega_{1}$ and a quickly-varying sinusoid with angular frequency $\omega_{2}.$ The low-frequency sinusoid is used here to describe slowly-varying (e.g., hourly) traffic demands, while the high-frequency sinusoid is used to model sudden (e.g., at the minute level) variations in traffic demand. For our simulations, we let $\theta_{0}=3$ veh/h, $\theta_{1}=1$ veh/h, $\theta_{2}=0.1$ veh/h, $\omega_{1}=0.1$ rad/hour, $\omega_{2}=\sqrt{50}$ rad/hour, and $\phi_{1}=\phi_{2}=0,$ see Fig. 6 (top).
We applied Algorithm 1 to derive an optimization algorithm to solve the traffic assignment problem (7.3). For the synthesis, we utilized the internal model $s(z)=Sz,$ where
$S=\text{diag}\biggl{(}\begin{bmatrix}0&1\\ -\omega_{1}^{2}&0\end{bmatrix},\begin{bmatrix}0&1\\ -\omega_{2}^{2}&0\end{bmatrix},0\biggr{)}.$
Notice that knowledge of $\theta_{0},\theta_{1},\theta_{2},\phi_{1},\phi_{2}$ is not required to synthesize the optimization algorithm – only the frequencies $\omega_{1},\omega_{2}$ are required to be known. To seek a solution to the constrained problem (7.3), consider the Lagrangian function: